A case-control (or retrospective) study was conducted to investigate a relationship between the colors of helmets worn by motIdentify the null and alternative hypotheses. Choose the correct answer below.

Compute the test statistic.

Find the critical​ value(s).

What is the conclusion based on the hypothesis​ test?
Answer #1

Please don't hesitate to give a "thumbs up" for the answer in case the answer has helped you

1.

Null hypothesis : Injuries are independent of helmet color

Alternate hypothesis: Injuries are NOT independent of helmet color

2.

The contingency table below provides the following information: the observed cell totals, (the expected cell totals) and [the chi-square statistic for each cell].

The chi-square statistic, p-value and statement of significance appear beneath the table. Blue means you're dealing with dependent variables; red, independent.

For example, for the 1st cell, it is:

O = 482

E = RowSum*ColumnSum/Total = 1121*696/1571 = 496.64

ChiSquare test statistic for 1st cell is: (O-E)^2/E = (482-496.64)^2/496.64 = 0.43

Results
black white yellow red blue Row Totals
control 482  (496.64)  [0.43] 350  (330.38)  [1.17] 32  (29.26)  [0.26] 167  (169.11)  [0.03] 90  (95.62)  [0.33] 1121
cases 214  (199.36)  [1.07] 113  (132.62)  [2.90] 9  (11.74)  [0.64] 70  (67.89)  [0.07] 44  (38.38)  [0.82] 450
Column Totals 696 463 41 237 134 1571  (Grand Total)

We add all cell' chi-squares and get 7.7173

Answer: The chi-square statistic is 7.7173.

Critical values are :

The p-value is .102499. The result is not significant at p < .05.

The chi-square critical statistic are: CHISQ.INV(0.025, 4) = 0.4844 and CHISQ.INV(0.975, 4) = 11.1433

Answer: Critical value(s): 0.4844, 11.1433

We fail to reject Ho as our p-value is more than .05.

Answer: Conclusion: We conclude that injuries are independent of helmet color.